OpenStudy (anonymous):

This point (-3, 4) is on the terminal side of an angle \theta in standard position. would \sin\theta=4?

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

Hmm this doesn't appear to be on the unit circle. when we're writing our coordinates in terms of sines and cosines it will look like, \( (r \cos \theta, r \sin \theta)\). So I think what we actually have is \(r \sin \theta=4\) And based on our picture we can determine the radius fairly easily using the Pythagorean Theorem.

zepdrix (zepdrix):

Sorry for the late response :C got busy.